OpenStudy (anonymous):

find a vector in the direction 2i+j-2k that has length 1

OpenStudy (amistre64):

how long is the given vector to begin with?

OpenStudy (amistre64):

vector = L units to determine the vector of 1 unit, divide both sides by L vector/L = L/L units vector/L = 1 units

OpenStudy (anonymous):

we just have the direction and the length of the required vector

OpenStudy (amistre64):

thats all we need, and some knowledge about what a unit vector is spose to be ....

OpenStudy (anonymous):

i have idea like: unit vector of that direction is <2, 1, -2>/sqr(2^2 +1^2+(-2)^2 =<2/sqr(5,1/sqr(5),-2/sqr(5)> is is right?

OpenStudy (amistre64):

thats correct :) except maybe check you length again

OpenStudy (anonymous):

and use the length 1 multiply to the unit vector to get what we need, is it right?

OpenStudy (amistre64):

sqrt(8) = 2sqrt(2)

OpenStudy (amistre64):

ugh ... 4+1+4 = 9 so sqrt(9) = 3

OpenStudy (anonymous):

yeah, you are right, my bad. thanks a lot

OpenStudy (amistre64):

divide all the vector parts by 3 :)